# SuperQ-GRASP: Superquadrics-based Grasp Pose Estimation on Larger Objects for Mobile-Manipulation

### v1

@@ Abstract
Grasp planning and estimation have been a long-standing research problem in robotics, with two main approaches to find graspable poses on the objects: 1) geometric approach, which relies on 3D models of objects and the gripper to estimate valid grasp poses, and 2) data-driven, learning-based approach, with models trained to identify grasp poses from raw sensor observations. The latter assumes comprehensive geometric coverage during the training phase. However, the data-driven approach is typically biased toward tabletop scenarios and struggle to generalize to out-of-distribution scenarios with larger objects (e.g. chair). Additionally, raw sensor data (e.g. RGB-D data) from a single view of these larger objects is often incomplete and necessitates additional observations. In this paper, we take a geometric approach, leveraging advancements in object modeling (e.g. NeRF) to build an implicit model by taking RGB images from views around the target object. This model enables the extraction of explicit mesh model while also capturing the visual appearance from novel viewpoints that is useful for perception tasks like object detection and pose estimation. We further decompose the NeRF-reconstructed 3D mesh into superquadrics (SQs) - parametric geometric primitives, each mapped to a set of precomputed grasp poses, allowing grasp composition on the target object based on these primitives. Our proposed pipeline overcomes the problems: a) noisy depth and incomplete view of the object, with a modeling step, and b) generalization to objects of any size. For more qualitative results, refer to the supplementary video and webpage https://bit.ly/3ZrOanU.
@@ I Introduction
The data-driven approach trains models to estimate grasp poses directly from raw sensor data, such as point clouds or images[8, 9, 10]. Although these models can handle outliers with sufficient training, they often suffer from biases in the training dataset. Most data-driven models are trained on tabletop scenarios, where objects are typically smaller, convex, or of low genus [10, 11, 8, 12, 13]. As a result, these methods struggle to generalize to larger, more complex objects and to accommodate limited observations across various viewpoints (see Fig. 1).
Consider manipulation tasks involving mobile manipulators that handle larger, relatively complex objects (concave and high genus) than those typically encountered in tabletop scenarios (convex and low genus). These tasks require grasping from diverse 3D viewpoints. While existing data-driven methods have proven effective in many applications, they face challenges when applied to such larger objects. They are typically trained on datasets containing smaller tabletop items, which may limit their generalization to more complex, larger objects.
@@ II-A Input Data Format for Robotic Grasping Task
Various input data formats have been employed to generate grasp poses for objects. Some approaches rely on visual input, particularly RGB-D images, as demonstrated in several works [18, 12, 19, 20, 21, 22, 23, 24]. For instance, [21, 25] adopt a multiview setup where images captured from multiple perspectives are processed by deep learning pipelines to predict grasp poses. In contrast, other studies [26, 27, 28, 29] focus on depth data in the form of point clouds for grasp pose prediction. For example, [26] utilizes a depth camera to capture point cloud data of the target object, which is then converted into a Truncated Signed Distance Function (TSDF) and fed into deep learning models.
Despite this, many existing methods operate in confined environments, such as tabletops, where data collection is limited to small spaces. For larger objects, as studied in this research, depth data acquisition is often incomplete and restricted to partial views, leading to inaccurate results (see Fig. 1).
@@ II-B Grasp Pose Estimation on a Target Object
Inspired by the works in [15, 38], we propose a non-deep learning approach to mitigate the bias inherent in training datasets used for data-driven grasp pose estimation methods. Unlike traditional geometric-based approaches, our method requires less precision in the object’s mesh, as it utilizes a technique robust to outliers. Moreover, by not relying on deep learning models, our approach offers an alternative that reduces dependence on large, curated datasets, enabling more flexible and unbiased grasp pose estimation.
@@ III Superquadrics Background
Figure 2: Grasp poses on individual superquadrics of different shapes, specified by different pairs of parameters ($\varepsilon_{1}$, $\varepsilon_{2}$) as grasp candidates
@@ V Methodology
Figure 3: Illustration of boundary points and locations of grasp poses in a quadrant on the individual superquadrics. For the continuous region, the grasp poses are obtained along the sample points on the mesh. For the discontinuous region, the grasp poses are sampled along the asymptotic lines.
@@ V-B Grasp Pose Sampling on SQs
We then connect a line between the boundary point and the edge of the cross-sectional plane for both boundary points. We sample grasp poses along the lines (See Fig. 3) using the same tolerance value as the one to generate valid grasp poses at continuous regions. Finally, we generate the grasp poses for the other quadrants using symmetry, and all grasp poses are combined as the grasp candidates $\{\hat{G}_{i}^{k}\}_{k=1}^{K}$ associated with this specific superquadric $SQ_{i}$.
@@ V-C Grasp Candidate Validation
In practice, to reduce computational cost, we iteratively select the closest superquadric to the current gripper pose until valid grasp poses are found. This iterative selection of superquadrics for grasp candidate extraction offers ways to determine the semantic portions of an object informed by a high-level task.
@@ VII-B Real-world Mobile Manipulation Experiment
Figure 6: Qualitative results of grasp poses on Chair 3 from synthetic data using the two baseline methods (CG+Mesh, CG+Depth) and our method (SuperQ-GRASP). Our method predicts more grasp poses than the baseline methods and provides grasp poses that are more concentrated in a specific region. Red indicates invalid grasps, while green indicates valid grasps
@@
Our pipeline demonstrates a higher success rate across four test objects (two chairs, a vacuum cleaner, and a table), highlighting its capability to estimate valid grasp poses for larger objects with complex geometries, including high-genus objects like chairs. However, in real-world trials, the primary source of failure when employing SuperQ-GRASP was due to pose estimation inaccuracies, which constrained the overall performance of both pipelines. For the luggage case scenario, both methods performed similarly, likely because the object’s limited graspable features—primarily the two handles—have a simple, stick-like structure that is easier to process. For more results, see our supplementary video and webpage: https://bit.ly/3ZrOanU.
@@ VIII Conclusion
This work is developed to facilitate the grasping of larger objects that are typically uncommon in tabletop scenarios for mobile manipulators. We present a comprehensive grasp generation pipeline that utilizes RGB images of large target objects to model the object and decompose it into primitive shapes for identifying valid grasp poses. We propose SuperQ-GRASP, a method that decomposes the object’s mesh into superquadrics and identifies valid grasps near the robot’s gripper. We validate the effectiveness of SuperQ-GRASP through experiments on our dataset, demonstrating its capability to generate valid grasp poses in proximity to the gripper. Additionally, we demonstrate the applicability of our comprehensive pipeline to tasks that require grasping large objects in mobile manipulation contexts. One limitation we recognize is the assumption of accurate information regarding the gripper’s pose relative to the object. Consequently, in real-world experiments, the performance of our grasp pose estimation method (SuperQ-GRASP) may be compromised when the pose estimation module fails to generate accurate results. Future studies can focus on enhancing the robustness of our method against inaccuracies in pose estimation results.
@@ Acknowledgement
The authors would like to thank Shirley Su and all other mem- bers of the Robotics: Perception and Manipulation (RPM) Lab for their insightful discussions and proofreading of this paper.

### v3

@@ Abstract
Grasp planning and estimation have been a long-standing research problem in robotics, with two main approaches to find graspable poses on the objects: 1) geometric approach, which relies on 3D models of objects and the gripper to estimate valid grasp poses, and 2) data-driven, learning-based approach, with models trained to identify grasp poses from raw sensor observations. The latter assumes comprehensive geometric coverage during the training phase. However, the data-driven approach is typically biased toward tabletop scenarios and struggle to generalize to out-of-distribution scenarios with larger objects (e.g. chair). Additionally, raw sensor data (e.g. RGB-D data) from a single view of these larger objects is often incomplete and necessitates additional observations. In this paper, we take a geometric approach, leveraging advancements in object modeling (e.g. NeRF) to build an implicit model by taking RGB images from views around the target object. This model enables the extraction of explicit mesh model while also capturing the visual appearance from novel viewpoints that is useful for perception tasks like object detection and pose estimation. We further decompose the NeRF-reconstructed 3D mesh into superquadrics (SQs) - parametric geometric primitives, each mapped to a set of precomputed grasp poses, allowing grasp composition on the target object based on these primitives. Our proposed pipeline overcomes the problems: a) noisy depth and incomplete view of the object, with a modeling step, and b) generalization to objects of any size. For more qualitative results, refer to the supplementary video and webpage https://rpm-lab-umn.github.io/superq-grasp-webpage/.
@@ I Introduction
The data-driven approach trains models to estimate grasp poses directly from raw sensor data, such as point clouds or images[8, 9, 10]. Although these models can handle outliers with sufficient training, they often suffer from biases in the training dataset. Most data-driven models are trained on tabletop scenarios, where objects are typically smaller, convex, or of low genus [10, 11, 8, 12, 13]. As a result, these methods struggle to generalize to larger, more complex objects of high genus and to accommodate limited observations across various viewpoints (see Fig. 1).
@@ II-A Input Data Format for Robotic Grasping Task
Various input data formats have been employed to generate grasp poses for objects. Some approaches rely on visual input, particularly RGB-D images, as demonstrated in several works [18, 12, 19, 20, 21, 22, 23, 24]. For instance, [21, 25] adopt a multiview setup where images captured from multiple perspectives are processed by deep learning pipelines to predict grasp poses. In contrast, other studies [26, 27, 28, 29] focus on depth data in the form of point clouds for grasp pose prediction. For example, [26] utilizes a depth camera to capture point cloud data of the target object, which is then converted into a Truncated Signed Distance Function (TSDF) and fed into deep learning models. However, these existing methods operate in confined environments, such as tabletops, where sensor observation is limited to small spaces. This limitation renders them less suitable for larger objects, as examined in this research, where depth data acquisition is frequently incomplete and constrained to partial views, resulting in reduced accuracy or even invalid grasp poses. (see Fig. 1).
@@ II-B Grasp Pose Estimation on a Target Object
Inspired by the works in [15, 38] that utilizes primitive decomposition to generate valid grasp poses, we propose a non-deep learning approach to mitigate the bias inherent in training datasets used for data-driven grasp pose estimation methods. Unlike traditional geometric-based approaches, our method requires less precision in the object’s mesh, as it utilizes a technique robust to outliers. Moreover, by not relying on deep learning models, our approach offers an alternative that reduces dependence on large, curated datasets, enabling more flexible and unbiased grasp pose estimation.
@@ III Superquadrics Background
Figure 2: Grasp pose candidates on individual superquadrics of different shapes, specified by different pairs of parameters ($\varepsilon_{1}$, $\varepsilon_{2}$)
@@ V Methodology
Figure 3: Illustration of boundary points and locations of grasp pose candidates in a quadrant on the individual superquadrics. For the continuous region, the grasp pose candidates are obtained along the sample points on the mesh. For the discontinuous region, the grasp pose candidates are sampled along the asymptotic lines.
@@ V-B Grasp Pose Sampling on SQs
We then connect an asymptotic line between the boundary point and the edge of the cross-sectional plane for both boundary points. We sample grasp poses along the lines (See Fig. 3) using the same tolerance value as the one to generate valid grasp poses at continuous regions. Finally, we generate the grasp poses for the other quadrants using symmetry, and all grasp poses are combined as the grasp candidates $\{\hat{G}_{i}^{k}\}_{k=1}^{K}$ associated with this specific superquadric $SQ_{i}$.
@@ V-C Grasp Candidate Validation
In practice, to reduce the computational cost, we iteratively select the closest superquadric to the current gripper pose until valid grasp poses are found. This iterative selection of superquadrics for grasp candidate extraction offers ways to determine the semantic portions of an object informed by a high-level task.
@@ VII-B Real-world Mobile Manipulation Experiment
Figure 6: Qualitative results of grasp poses on Chair 3 from the synthetic dataset using the two baseline methods (CG+Mesh, CG+Depth) and our method (SuperQ-GRASP). Our method predicts more grasp poses than the baseline methods and provides grasp poses that are more concentrated in a specific region. Red indicates invalid grasps, while green indicates valid grasps
@@
Our pipeline demonstrates a higher success rate across four test objects (two chairs, a vacuum cleaner, and a table), highlighting its capability to estimate valid grasp poses for larger objects with complex geometries, including high-genus objects like chairs. However, in real-world trials, the primary source of failure when employing SuperQ-GRASP was due to pose estimation inaccuracies, which constrained the overall performance of both pipelines. For the luggage case scenario, both methods performed similarly, likely because the object’s limited graspable features—primarily the two handles—have a simple, stick-like structure that is easier to process. For more results, including the execution time, see our supplementary video and webpage: https://rpm-lab-umn.github.io/superq-grasp-webpage/.
@@ VIII Conclusion
In this work, we propose SuperQ-GRASP, a method that decomposes the object’s mesh into superquadrics and identifies valid grasps near the robot’s gripper. Based on the proposed method, we also present a comprehensive grasp generation pipeline that utilizes RGB images of large target objects uncommon in tabletop scenarios to model the object and identify valid grasp poses. We validate the effectiveness of SuperQ-GRASP through experiments on our dataset, demonstrating its capability to generate valid grasp poses in proximity to the gripper. Additionally, we demonstrate the applicability of our comprehensive pipeline to tasks that require grasping large objects in mobile manipulation contexts. One limitation we recognize is the assumption of accurate information regarding the gripper’s pose relative to the object. Consequently, in real-world experiments, the performance of our grasp pose estimation method (SuperQ-GRASP) may be compromised when the pose estimation module fails to generate accurate results. Future studies can focus on enhancing the robustness of our method against inaccuracies in pose estimation results.
@@ Acknowledgement
The authors would like to thank Guanang (Shirley) Su and all other members of the Robotics: Perception and Manipulation (RPM) Lab for their insightful discussions and proofreading of this paper.